Physics HW question. (I want you to show me all the work)

*I want you to show me all the work*

An charge with mass m and charge q is emitted from the origin,(x,y)=(0,0) . A large, flat screen is located at x = L . There is a target on the screen at y position yh, where yh>0 . In this problem, you will examine two different ways that the charge might hit the target. Ignore gravity in this problem.

a.Assume that the charge is emitted with velocity v0 in the positive x direction. Between the origin and the screen, the charge travels through a constant electricfield pointing in the positive y direction. What should the magnitude E of the electric field be if the charge is to hit the target on the screen?
(Express your answer in terms of m, q, yh, v0, and L)
E = ____________

b.Now assume that the charge is emitted with velocity v0 in the positive y direction. Between the origin and the screen, the charge travels through a constant electricfield pointing in the positive x direction. What should the magnitude of E the electric field be if the charge is to hit the target on the screen?
(Express your answer in terms of m, q, yh, v0, and L)
E = ____________

*I want you to show me all the work*
Answer #1

part a:

Break it down into the motions of the x and y directions.

Motion in X-direction:

x(t) = x(0) + (v0)t

Therefore: L = 0 + (v0)t
t = (L/v0)

Motion in Y-direction:

y(t) = y(0) = .5at^2

remember a = qE/m

y_h = y(0) + .5(qE/m)(L/v0)^2

Now, just solve for E.

E = (2*m*y_h*(v0)^2) / (q*L^2)

For Part B:

Simply reverse the equation so that it has negative values instead of positive.
Answer #2

This video explains it! https://youtu.be/AGGprvExV8g 

answered by: Super Cool School